From a box containing 4 defective and 5 non defective items,



how many samples of size 3 are possible?

1
Expert's answer
2021-11-23T08:49:17-0500

The box contains both defective and non defective items. Let "D" be the event that an item is defective and "D'" be the event that an item is non defective. Therefore, there are "k=2" sets of events. If samples of size "n=3" are drawn from this box then, we expect "k^n" possible samples. Therefore, the number of possible samples are, "k^n=2^3=8".